# An implicit–explicit BDF–Galerkin scheme of second order for the nonlinear thermistor problem

### v1

@@ 6.2. Extension to higher order
This subsection is devoted to a numerical study of an implicit-explicit BDF–Galerkin scheme of third order in time. Following the construction of Section 2, we combine the standard BDF-$3$ operator
@@
To initialize the method for $n\leq 2$, we use the exact solution. For $n=3,\ldots,N$, the fully discrete third-order BDF–Galerkin FEM then takes the form
@@
Since we expect an $L^{2}$-error of order $\mathcal{O}(h^{2}+h\tau^{3})$, we choose $\tau=h^{1/3}$ for the time step size. Indeed, this leads to a convergence rate of order two. Also for the other error components, we obtain the expected optimal rates, i.e., order one in the $H^{1}$-norm and order two in the supercloseness and superconvergence estimates. To verify the expected third-order in time, we fix a mesh size $h$ and reduce $\tau$; see Figure 6.6.
@@ 6.3. Comparison to related methods
In this final subsection, we highlight the advantages of our method and present a comparison with related numerical schemes. First, we consider the second-order BDF–Galerkin FEM investigated in [Gao16], which reads
@@
for all $\xi_{u},\,\xi_{\phi}\in V_{h,0}$. Note that, compared to scheme (2.11), different extrapolation formulae are used in both equations. To compare the results with the previous experiments, we choose again $\tau=\sqrt{h}$. As shown in Table 6.1, this reduces the convergence order in the $L^{2}$-error as well as in the interpolation-type errors. If we set $\tau=h$ in the same experiment, we regain second-order convergence in the $L^{2}$-error and superconvergence. Supercloseness of order two, however, is not reached.
Table 6.1. Numerical errors and convergence orders in $u$ and $\phi$ at time $t=T$ using the BDF–Galerkin FEM (6.2) and $\tau=\sqrt{h}$.
mesh $8\times 8$ $16\times 16$ $32\times 32$ $64\times 64$
$\|U_{h}^{N}-u^{N}\|_{0}$ $2.016\cdot 10^{-2}$ $1.559\cdot 10^{-2}$ $8.947\cdot 10^{-3}$ $4.328\cdot 10^{-3}$
order $0.37$ $0.80$ $1.05$
$\|U_{h}^{N}-I_{h}u^{N}\|_{1}$ $1.458\cdot 10^{-1}$ $8.781\cdot 10^{-2}$ $4.411\cdot 10^{-2}$ $2.133\cdot 10^{-2}$
order $0.73$ $0.99$ $1.05$
$\|I_{2h}U_{h}^{N}-u^{N}\|_{1}$ $1.018\cdot 10^{-1}$ $8.240\cdot 10^{-2}$ $4.507\cdot 10^{-2}$ $2.099\cdot 10^{-2}$
order $0.31$ $0.87$ $1.10$
$\|\Phi_{h}^{N}-\phi^{N}\|_{1}$ $7.789\cdot 10^{-2}$ $3.957\cdot 10^{-2}$ $1.935\cdot 10^{-2}$ $9.765\cdot 10^{-3}$
order $0.98$ $1.03$ $0.99$
$\|\Phi_{h}^{N}-I_{h}\phi^{N}\|_{1}$ $7.856\cdot 10^{-2}$ $3.973\cdot 10^{-2}$ $1.941\cdot 10^{-2}$ $9.778\cdot 10^{-3}$
order $0.98$ $1.03$ $0.99$
$\|I_{2h}\Phi_{h}^{N}-\phi^{N}\|_{1}$ $1.307\cdot 10^{-2}$ $2.807\cdot 10^{-3}$ $7.290\cdot 10^{-4}$ $2.208\cdot 10^{-4}$
order $2.22$ $1.95$ $1.72$
@@
for all $\xi_{u},\,\xi_{\phi}\in V_{h,0}$. Using again $\tau=\sqrt{h}$, the convergence, supercloseness, and superconvergence results are reported in Table 6.2. From this table, we observe that the $L^{2}$-error as well as the interpolation-type errors show irregular behavior on coarse meshes and overall exhibit a drop of the convergence rates.
@@
Table 6.2. Numerical errors and convergence orders in $u$ and $\phi$ at time $t=T$ using the BDF–Galerkin FEM (6.3) and $\tau=\sqrt{h}$.
mesh $8\times 8$ $16\times 16$ $32\times 32$ $64\times 64$
$\|U_{h}^{N}-u^{N}\|_{0}$ $8.168\cdot 10^{-4}$ $1.503\cdot 10^{-3}$ $8.062\cdot 10^{-4}$ $3.051\cdot 10^{-4}$
order 0.88-0.88 $0.90$ $1.40$
$\|U_{h}^{N}-I_{h}u^{N}\|_{1}$ $4.753\cdot 10^{-3}$ $7.223\cdot 10^{-3}$ $3.727\cdot 10^{-3}$ $1.394\cdot 10^{-3}$
order 0.60-0.60 $0.96$ $1.42$
$\|I_{2h}U_{h}^{N}-u^{N}\|_{1}$ $7.171\cdot 10^{-2}$ $1.812\cdot 10^{-2}$ $5.464\cdot 10^{-3}$ $1.895\cdot 10^{-3}$
order $1.98$ $1.73$ $1.53$
$\|\Phi_{h}^{N}-\phi^{N}\|_{0}$ $1.025\cdot 10^{-3}$ $5.430\cdot 10^{-4}$ $2.251\cdot 10^{-4}$ $1.547\cdot 10^{-4}$
order $0.92$ $1.27$ $0.54$
$\|\Phi_{h}^{N}-I_{h}\phi^{N}\|_{1}$ $6.757\cdot 10^{-3}$ $3.541\cdot 10^{-3}$ $1.512\cdot 10^{-3}$ $1.025\cdot 10^{-3}$
order $0.93$ $1.23$ $0.56$
$\|I_{2h}\Phi_{h}^{N}-\phi^{N}\|_{1}$ $1.251\cdot 10^{-2}$ $3.379\cdot 10^{-3}$ $1.438\cdot 10^{-3}$ $1.002\cdot 10^{-3}$
order $1.89$ $1.23$ $0.52$
@@ 7. Concluding Remarks
While we have focused on BDF of second order, the analytical techniques developed here may be extended to higher-order methods with appropriate extrapolation terms. In particular, this includes the temporal–spatial error splitting technique and the telescope formula of the temporal discretization operator. Moreover, the techniques presented in this work may be extended to derive implicit–explicit methods for other nonlinear systems.
Acknowledgments

### v2

@@ 6.2. Extension to higher order
This subsection is devoted to a numerical study of an implicit–explicit BDF–Galerkin scheme of third order in time. Following the construction of Section 2, we combine the standard BDF-$3$ operator
@@
To initialize the method for $n\leq 2$, we use the exact solution, ensuring that the initial error does not affect the third-order accuracy. Alternatively, one may follow [GAO16] and compute the starting values by a lower-order initialization procedure involving an implicit–explicit Euler step and Crank–Nicolson-type corrections. This then yields approximations with temporal error of order $O(\tau^{3})$. Such a construction preserves the global third-order accuracy of the scheme. For $n=3,\ldots,N$, the fully discrete third-order BDF–Galerkin FEM then takes the form
@@
Since we expect an $L^{2}$-error of order $\mathcal{O}(h^{2}+\tau^{3})$, we choose $\tau=h^{2/3}$ for the time step size. Indeed, this leads to a convergence rate of order two. Also for the other error components, we obtain the expected optimal rates, i.e., order one in the $H^{1}$-norm and order two in the supercloseness and superconvergence estimates. Finally, to verify the expected third-order convergence in time, we fix a sufficiently fine spatial mesh and reduce the time step size. The resulting $L^{2}$-errors at time $t=T$ are plotted in Figure 6.3, indicating third-order convergence in time.
@@ 6.3. Comparison to related methods
In this subsection, we highlight the advantages of our method and present a comparison with related numerical schemes. First, we consider the second-order BDF–Galerkin FEM investigated in [GAO16], which reads
@@
for all $\xi_{u},\,\xi_{\phi}\in V_{h,0}$. Note that, compared to scheme (2.11), different extrapolation formulae are used in both equations. To compare the results with the previous experiments, we choose again $\tau=h$. As shown in Table 6.3, this reduces the convergence order in the interpolation-type errors.
Table 6.3. Numerical errors and convergence orders in $u$ and $\phi$ at time $t=T$ using the BDF–Galerkin FEM (6.3) with $\tau=h$.
mesh $32\times 32$ $64\times 64$ $128\times 128$ $256\times 256$
$\|U_{h}^{N}-I_{h}u^{N}\|_{1}$ $2.17\cdot 10^{-2}$ $1.04\cdot 10^{-2}$ $5.25\cdot 10^{-3}$ $2.62\cdot 10^{-3}$
order $1.05$ $0.99$ $1.00$
$\|\Phi_{h}^{N}-I_{h}\phi^{N}\|_{1}$ $2.70\cdot 10^{-2}$ $1.34\cdot 10^{-2}$ $6.69\cdot 10^{-3}$ $3.35\cdot 10^{-3}$
order $1.01$ $1.00$ $1.00$
@@
for all $\xi_{u},\,\xi_{\phi}\in V_{h,0}$. Using again $\tau=h$, not only the interpolation-type error $\|\Phi_{h}^{N}-I_{h}\phi^{N}\|_{1}$ but also the $L^{2}$-error $\|\Phi_{h}^{N}-\phi^{N}\|_{0}$ shows a reduced convergence order of one.
6.4. Further experiments
To demonstrate the validity of the theoretical analysis for more realistic thermistor problems, we study a number of additional experiments. In the following, we only discuss the numerical results for $u$, since the approximations of $\phi$ exhibit the corresponding convergence behavior.
6.4.1. Three-dimensional problem
Although not covered by the theoretical analysis, we also consider a three-dimensional numerical experiment. For this, we consider problem (6.1) on the unit cube $\Omega=(0,1)^{3}$ with right-hand sides such that the exact solution is given by
$u(x,y,z,t)=e^{-2t}\sin(\pi x)\sin(\pi y)\sin(\pi z),\qquad\phi(x,y,z,t)=1+\sin(x+y+z+t).$
We partition the domain into cubes with grid resolutions of $M\times M\times M$ with $M=4,8,16$. The corresponding mesh size is given by $h=\sqrt{3}/M$ and we set $\tau=h$ for the time step size. The convergence, supercloseness, and superconvergence results of the proposed BDF–Galerkin scheme (2.11) for $u$ at time $t=T$ are reported in Table 6.4. One can see that the expected convergence behavior is retained also in three space dimensions.
@@ 6.4.1. Three-dimensional problem
Table 6.4. Numerical errors for $u$ at final time $t=T$ using the second-order BDF–Galerkin FEM (2.11) for the three-dimensional problem.
mesh $4\times 4\times 4$ $8\times 8\times 8$ $16\times 16\times 16$
$\|U_{h}^{n}-u^{n}\|_{0}$ $2.4449\cdot 10^{-3}$ $5.5167\cdot 10^{-4}$ $1.2898\cdot 10^{-4}$
order $2.15$ $2.10$
$\|U_{h}^{n}-u^{n}\|_{1}$ $7.2627\cdot 10^{-2}$ $3.2521\cdot 10^{-2}$ $1.7377\cdot 10^{-2}$
order $1.16$ $0.90$
$\|U_{h}^{n}-I_{h}u^{n}\|_{1}$ $2.5875\cdot 10^{-2}$ $5.1381\cdot 10^{-3}$ $1.0929\cdot 10^{-3}$
order $2.33$ $2.23$
$\|I_{2h}U_{h}^{n}-u^{n}\|_{1}$ $7.8195\cdot 10^{-2}$ $1.1407\cdot 10^{-2}$ $3.7514\cdot 10^{-3}$
order $2.78$ $1.60$
6.4.2. Different conductivities
In the upcoming examples, we consider the original system (1.1) without artificial forcing. To evaluate the spatial convergence behavior, a reference solution is computed on a fine mesh ($M=256$, $\tau=2^{-9}$). The computations are then performed on uniform rectangular meshes with mesh size $h=\sqrt{2}/M$, $M=16,32,64$, and time step size $\tau=h$.
We investigate the numerical performance for different representative choices of the conductivity function $\sigma$. For this, we consider a strongly decreasing
(6.5) $\sigma(u)=\frac{2}{1+e^{u}}+0.1$
as well as an increasing behavior,
(6.6) $\sigma(u)=0.5+0.5\tanh(u).$
Such nonlinear conductivities are admissible under the classical thermistor theory, provided they are continuous, uniformly positive, and bounded [AC94]. Similar choices are standard in both analytical and numerical studies of thermistor problems [CF93, DA19]. Convergence rates in the $L^{2}$- and $H^{1}$-norms, as well as superconvergence-related quantities, are evaluated in order to assess whether the theoretical error estimates suffer from different nonlinear conductivity laws. The resulting errors for the strongly decreasing conductivity (6.5) are listed in Table 6.5, indicating that the proposed method preserves its optimal convergence and superconvergence properties. Very similar results are obtained for the increasing conductivity (6.6) such that we omit the actual convergence results.
Table 6.5. Numerical errors for the temperature $u$ at final time $t=T$ for the thermistor problem with strongly decreasing conductivity (6.5), computed by the second-order BDF–Galerkin FEM (2.11) with $\tau=h$.
mesh $16\times 16$ $32\times 32$ $64\times 64$
$\|U_{h}^{n}-u^{n}\|_{0}$ $2.710\times 10^{-4}$ $6.812\times 10^{-5}$ $1.720\times 10^{-5}$
order $1.99$ $1.99$
$\|U_{h}^{n}-u^{n}\|_{1}$ $2.234\times 10^{-2}$ $1.100\times 10^{-2}$ $4.855\times 10^{-3}$
order $1.02$ $1.18$
$\|U_{h}^{n}-I_{h}u^{n}\|_{1}$ $2.163\times 10^{-3}$ $6.389\times 10^{-4}$ $1.812\times 10^{-4}$
order $1.76$ $1.82$
$\|I_{2h}U_{h}^{n}-u^{n}\|_{1}$ $9.252\times 10^{-4}$ $2.401\times 10^{-4}$ $6.031\times 10^{-5}$
order $1.95$ $1.99$
6.4.3. Non-convex domains
Next, we consider system (1.1) on the non-convex L-shape domain
$\Omega=(0,1)^{2}\,/\,([0.5,1]\times[0.5,1]),$
which is known to cause reduced regularity in elliptic and parabolic problems. In this case, the boundary conditions are selected to generate nontrivial electric currents, i.e., $g(x,y)=x$, while no artificial source terms are added to the governing equations. This example allows us to examine the influence of geometric singularities on the numerical solution. The resulting convergence behavior for $u$, showing degraded rates in all measures, is given in Table 6.6.
Table 6.6. Numerical errors for $u$ at final time $t=T$ on the L-shape, computed with the second-order BDF–Galerkin FEM (2.11) and $\tau=h$.
mesh $16\times 16$ $32\times 32$ $64\times 64$
$\|U_{h}^{n}-u^{n}\|_{0}$ $7.709\cdot 10^{-4}$ $2.199\cdot 10^{-4}$ $6.657\cdot 10^{-5}$
order $1.81$ $1.72$
$\|U_{h}^{n}-u^{n}\|_{1}$ $4.418\cdot 10^{-2}$ $2.416\cdot 10^{-2}$ $1.358\cdot 10^{-2}$
order $0.87$ $0.83$
$\|U_{h}^{n}-I_{h}u^{n}\|_{1}$ $1.194\cdot 10^{-4}$ $5.204\cdot 10^{-5}$ $2.033\cdot 10^{-5}$
order $1.20$ $1.36$
$\|I_{2h}U_{h}^{n}-u^{n}\|_{1}$ $7.763\cdot 10^{-4}$ $2.225\cdot 10^{-4}$ $6.533\cdot 10^{-5}$
order $1.80$ $1.77$
6.4.4. Non-smooth boundary data
Finally, we investigate the performance of the method under reduced regularity due to non-smooth boundary data. For this, the temperature satisfies homogeneous Dirichlet conditions as before, whereas the electric potential is equipped with discontinuous boundary conditions of the form $g(x,y)=0$ for $x<0.5$ and $g(x,y)=1$ for $x\geq 0.5$. The corresponding errors are reported in Table 6.7. As expected, the discontinuity in the boundary data leads to reduced regularity of the solution and a dramatic drop of the convergence rates. While the $L^{2}$-errors still show convergence of order one, convergence in the $H^{1}$-norm is noticeably slower.
Table 6.7. Numerical errors for $u$ at final time $t=T$ with non-smooth boundary conditions, computed using the second-order BDF–Galerkin FEM (2.11) and $\tau=h$.
mesh $16\times 16$ $32\times 32$ $64\times 64$
$\|U_{h}^{n}-u^{n}\|_{0}$ $3.438\times 10^{-2}$ $1.767\times 10^{-2}$ $8.251\times 10^{-3}$
order $0.96$ $1.10$
$\|U_{h}^{n}-u^{n}\|_{1}$ $1.596\times 10^{0}$ $1.360\times 10^{0}$ $1.054\times 10^{0}$
order $0.23$ $0.37$
$\|U_{h}^{n}-I_{h}u^{n}\|_{1}$ $1.582\times 10^{0}$ $1.334\times 10^{0}$ $1.012\times 10^{0}$
order $0.25$ $0.40$
$\|I_{2h}U_{h}^{n}-u^{n}\|_{1}$ $2.719\times 10^{-1}$ $2.492\times 10^{-1}$ $2.095\times 10^{-1}$
order $0.13$ $0.25$
@@ 7. Concluding Remarks
While we have focused on BDF of second order, the analytical techniques developed here may be extended to higher-order methods with appropriate extrapolation terms. In particular, this includes the temporal–spatial error splitting technique and the telescope formula of the temporal discretization operator. Also the convergence analysis of the three-dimensional case is a promising field of future research. Moreover, the techniques presented in this work may be extended to derive implicit–explicit methods for other nonlinear systems.